# Preempting Fermion Sign Problem: Unveiling Quantum Criticality through Nonequilibrium Dynamics

### v1

@@ IV.3 Phase boundary and critical exponents
IV.4 The new universality class
Our numerical results demonstrate the presence of a continuous phase transition which does not belong to any of known universality classes [71, 72]. Instead, it represents a new universality class.
@@ IV.4 The new universality class
In the saturated $\lambda_{8}$-AFM phase, two flavors of fermions (referred to as flavors 1 and 2) are localized in one sublattice (designated as the A lattice), while the remaining flavor of fermions (referred to as flavor 3) is localized in the other sublattice (designated as the B lattice), as illustrated in Fig. 3a of the main text. It is evident that the $\mathrm{Z}_{2}$ symmetry between the AB sublattices is significantly broken. The diagonal $\lambda_{8}$ generator induces a global U(1) transformation, while the generators $\lambda_{1},\lambda_{2},\lambda_{3}$ act solely on the subspace of flavors 1 and 2, generating a closed SU(2) transformation that only mixes these two flavors. Since these four generators all commute with the order parameter operator, the $\mathrm{SU(2)}\times\mathrm{U(1)}$ is the largest symmetry group that remains invariant under the $\lambda_{8}$-AFM order parameter $m$ after spontaneous symmetry breaking. The other four generators of the SU(3) group, $\lambda_{4},\lambda_{5},\lambda_{6},\lambda_{7}$, produce transformations that alter the direction of the AFM order parameter $\left\{m_{n}\right\}$ in a compact manifold, leading to other $\lambda_{8}$-AFM degenerate ground states, which are four independent Goldstone modes. All degenerate ground states span a 4-dimensional ground state degeneracy space, and these states can be mapped one-to-one onto points on the manifold $\left[\mathrm{SU(3)}\times\mathrm{Z}_{2}\right]/\left[\mathrm{SU(2)}\times\mathrm{U(1)}\right]$ [73, 74].
The critical exponents determined via the short-time dynamics in our work are summarized in Table S4 and compared with those of other Gross-Neveu universality classes with the same fermion components. It is clear that the new universality class we have discovered is distinct from the previously known classes.
@@
Table S4: Comparison of critical exponents for different Gross-Neveu-Yukawa universality classes in $d=2+1$ with $N_{\mathrm{f}}=6$. The first row, chiral $[\mathrm{SU(3)}\times\mathrm{Z_{2}}]/[\mathrm{SU(2)}\times\mathrm{U(1)}]$ denotes the universality class for $SU(3)$ Dirac fermion Hubbard model, with exponents determined from nonequilibrium PQMC.
@@
chiral $\mathrm{[SU(3)\times Z_{2}}]/\left[\mathrm{SU(2)}\times\mathrm{U(1)}\right]$ (this work) 0.68(5) 0.55(5) 0.15(3)
@@
IV.5 Sign problem behaviors and computational efficiency
@@
For this model, we set the parameters at the critical point $U=2.0(3)$, $\phi=0.1\pi$. The average sign as a function of imaginary time $\tau$ and size $L$ is shown in Fig. S13a. The sign problem for this model is significantly more severe than for the previous two models. For $\tau=0.25L^{z}$, the average sign is approximately $10^{-1}\sim 10^{-2}$, meaning that compared to the sign-problem-free case, $10$ to $100$ times more computational resources are needed to obtain reliable results. From Figs. S11 and S12 and Fig. 3 in the main text, it can be seen that $\tau=0.25L^{z}$ is in the nonequilibrium scaling region controlled by the critical point. Even though the ground state is not reached, its nonequilibrium scaling still reflects the quantum criticality of the ground state. In equilibrium QMC studies, it is typically necessary to set the imaginary time $\tau$ to more than $1.5$ times the size $L^{z}$ to reach the ground state. However, evolving for such a long time, the average sign decays to approximately $10^{-5}\sim 10^{-6}$. In Fig. S13b, we compare the average sign for $\tau=0.25L^{z}$ and $\tau=1.5L^{z}$. Since the computational error (according to Eq. S8) is inversely proportional to the average sign $\braket{\mathrm{sign}}$ (according to Eq. S8), we can measure the difference in computational efficiency by multiplying the ratio of the average sign $\braket{\mathrm{sign}}$ by the length of imaginary-time evolution, as follows:
$\mathrm{Efficiency~{}acceleration~{}factor}=\frac{{1/\braket{\mathrm{sign}}_{\mathrm{eq.}}}}{{1/\braket{\mathrm{sign}}_{\mathrm{neq.}}}}\times\frac{\tau_{\mathrm{eq.}}}{\tau_{\mathrm{neq.}}},$ (S31)

### v2

@@ DISCUSSION
S4.4 The new universality class
In three-dimensional classical systems or 2+1 dimensional quantum systems, the most typical O($N$) universality classes widely describe a variety of important phase transitions, such as the gas-liquid transition of simple gases, the superfluid transition of liquid helium, and the Heisenberg transition of ferromagnets. Specifically, the $N=1$ case is referred to as the Ising universality class, $N=2$ as the XY universality class, and $N=3$ as the Heisenberg universality class. Their critical behaviors can be simply described by purely bosonic $\phi^{4}$ theories with O($N$) symmetries.
When the boson order parameter fields are coupled to Dirac fermions, the critical properties are significantly modified by the gapless fermion fluctuations, which leads to the definition of the “chiral versions” of the above universality classes [74], namely the O($N_{\text{b}}$) universality classes. To avoid confusion, we here use $N$ to denote the dimension of the classical order parameter, $N_{\text{b}}$ to denote the dimension of the bosonic order parameter, and $N_{\text{f}}$ to denote the number of fermion flavors. For example, quantum phase transitions with a $Z_{2}$ charge-density-wave order parameter belong to the chiral Ising universality class; those with an O(2), or equivalently U(1), superconducting order parameter belong to the chiral XY universality class; and those with an O(3) antiferromagnetic order parameter belong to the chiral Heisenberg universality class. The corresponding low-energy effective field theories at criticality are obtained by coupling the $\phi^{4}$ bosonic theory of the order parameter to chiral fermion fields through Yukawa terms, and can generally be written as
$\mathcal{L}=\text{tr}\bigl(\bar{\psi}\not{\partial}\psi\bigr)+g\,\text{tr}\bigl(\bar{\psi}\Phi\psi\bigr)+\text{tr}\bigl((\partial_{\mu}\Phi)^{2}\bigr)+r\,\text{tr}\bigl(\Phi^{2}\bigr)+\lambda\,\bigl(\text{tr}(\Phi^{2})\bigr)^{2}\,$, (S35)
where $\psi$ is the fermion spinor field, and tr denotes the trace over the $N_{\text{f}}$ flavors of chiral fermions with SU($N_{\text{f}}$) symmetry. The operator $\Phi=\sum_{i=1}^{N_{\text{b}}}\phi_{i}L_{i}$ spans an $N_{\text{b}}$-dimensional linear representation of the SU($N_{\text{f}}$) algebra, with $N_{\text{b}}$ linearly independent bosonic components, and the $N_{\text{f}}$-dimensional matrices $L_{i}$ form an orthogonal basis of this representation, such that $\text{tr}(\Phi^{2})=\sum_{i=1}^{N_{\text{b}}}\phi_{i}^{2}$. This theory is the well-known Gross-Neveu-Yukawa theory. Therefore, these chiral O($N_{\text{b}}$) universality classes are collectively referred to as the Gross-Neveu universality classes. Their order parameters correspond to the $N_{\text{b}}$-dimensional real linear representation of the fermionic SU(2) algebra, which happens to carry a faithful representation of the O($N_{\text{b}}$) group. Consequently, the bosonic fields in the theory also possess O($N_{\text{b}}$) symmetry, establishing a one-to-one correspondence with the usual O($N$) universality classes. For instance, for the $N_{\text{b}}=3$ chiral Heisenberg universality class in an SU($N_{\text{f}}=2$) system, Eq. (S35) can be written in the more familiar form
$\mathcal{L}=\bar{\psi}\not{\partial}\psi+g\,\bar{\psi}\left(\vec{\phi}\cdot\vec{\sigma}\right)\psi+\left|\partial_{\mu}\vec{\phi}\right|^{2}+r\,\left|\vec{\phi}\right|^{2}+\lambda\,\left|\vec{\phi}\right|^{4}\,$, (S36)
where $\vec{\phi}=(\phi_{1},\phi_{2},\phi_{3})$ is an O(3) vector and $\vec{\sigma}=(\sigma_{x},\sigma_{y},\sigma_{z})$ are the Pauli matrices, forming an orthogonal basis for the three-dimensional linear representation of the SU(2) algebra. We summarize in Table S4 the theoretical symmetries and the vacuum manifolds (ground-state degeneracy manifolds) after spontaneous symmetry breaking corresponding to various chiral O($N_{\text{b}}$) universality classes.
Table S4: Comparison of the symmetries of different Gross-Neveu universality classes. The symmetries of the theories and the vacuum manifolds (ground-state degeneracy manifolds) after spontaneous symmetry breaking are listed. Each chiral O($N_{\text{b}}$) universality class corresponds to an O($N$) universality class with the same symmetries and vacuum manifold. In contrast, the new chiral SU(3) universality class reported in our manuscript has no corresponding O($N$) universality class.
Gross-Neveu classes Symmetry Vacuum manifold
chiral O($N_{\text{b}}$) $\left\{\begin{array}[]{l}\text{chiral Ising}\\ \text{chiral XY}\\ \text{chiral Heisenberg}\end{array}\right$. O($N_{\text{b}}$) $\left\{\begin{array}[]{l}\text{$Z_{2}$, $N_{\text{b}}=1$ }\\ \text{$O(2)$, $N_{\text{b}}=2$}\\ \text{$O(3)$, $N_{\text{b}}=3$}\end{array}\right$. S0
S1
S2
   chiral SU(3) $\text{SU(3)}\times\text{Z}_{2}$ $\frac{\text{SU(3)}\times\text{Z}_{2}}{\text{SU(2)}\times\text{U(1)}}\left(\simeq\mathbb{CP}^{2}\times\text{Z}_{2}\right)$
The Gross-Neveu universality class describes the universal physics in which massless Dirac fermions acquire a mass through spontaneous symmetry breaking. Although the known Gross-Neveu universality classes possess an O($N$) order-parameter structure, considering the intrinsic spinor nature of fermions, the Gross-Neveu universality class should not simply be regarded as the chiral counterpart of the classical O($N$) universality class. A fundamental question of broad interest across statistical physics, condensed matter physics, and high-energy physics is whether fermionic Gross-Neveu criticality can arise that goes beyond the O($N$) counterparts.
We find that SU($N_{\text{f}}$) Dirac fermion systems with with odd $N_{\text{f}}$ offer precisely such a pathway. From the perspective of the structure of symmetry group algebras, the key distinction between odd and even $N_{\text{f}}$ lies in the fact that SU($N_{\text{f}}$) groups with odd $N_{\text{f}}$ cannot be isomorphic to spin groups, which may allow for the emergence of “more fermionic” universality classes, as illustrated in Fig S17. For even $N_{\text{f}}$, however, at small values of $N_{\text{f}}$, such as SU(2) and SU(4), the groups are locally isomorphic to O(3) and O(6), respectively; while at large $N_{\text{f}}$, the system typically does not favor breaking SU($N_{\text{f}}$) symmetry to form spin order, but instead tends to form a valence-bond solid [82, 83]. Indeed, all Gross-Neveu transitions studied so far for even $N_{\text{f}}$ fall back into the conventional chiral O($N_{\text{b}}$) universality classes [84, 85, 86, 87, 88, 89]. Therefore, investigating the case of odd $N_{\text{f}}$ is of fundamental importance for understanding fermionic phase transitions.
Figure S17: For odd $N_{\text{f}}$, the SU($N_{\text{f}}$) group cannot be isomorphic to a spin group, which may give rise to new Gross-Neveu universality classes that more directly manifest the spinor nature of fermions.
We clearly identify the first nontrivial fermionic Gross-Neveu criticality that goes beyond the O($N$) order-parameter structures. The continuous transition from the SU(3) DSM to the $\lambda_{8}$-AFM phase reported in our manuscript belongs to the Gross-Neveu universality class family but does not fall into any of the known chiral Ising, chiral XY, or chiral Heisenberg classes.
@@
In the saturated $\lambda_{8}$-AFM phase, two flavors of fermions (referred to as flavors 1 and 2) are localized in one sublattice (designated as the A lattice), while the remaining flavor of fermions (referred to as flavor 3) is localized in the other sublattice (designated as the B lattice), as illustrated in Fig. 3a of the main text. It is evident that the $\mathrm{Z}_{2}$ symmetry between the AB sublattices is significantly broken. The diagonal $\lambda_{8}$ generator induces a global U(1) transformation, while the generators $\lambda_{1},\lambda_{2},\lambda_{3}$ act solely on the subspace of flavors 1 and 2, generating a closed SU(2) transformation that only mixes these two flavors. Since these four generators all commute with the order parameter operator, the $\mathrm{SU(2)}\times\mathrm{U(1)}$ is the largest symmetry group that remains invariant under the $\lambda_{8}$-AFM order parameter after spontaneous symmetry breaking. The other four generators of the SU(3) group, $\lambda_{4},\lambda_{5},\lambda_{6},\lambda_{7}$, produce transformations that alter the direction of the AFM order parameter $\left\{m_{n}\right\}$ in a compact manifold, leading to other $\lambda_{8}$-AFM degenerate ground states, which are four independent Goldstone modes. All degenerate ground states span a 4-dimensional ground state degeneracy space, and these states can be mapped one-to-one onto points on the manifold $\frac{\text{SU(3)}\times\text{Z}_{2}}{\text{SU(2)}\times\text{U(1)}}$. Under SU(3) transformations, this order parameter transforms according to the 8-dimensional adjoint representation of SU(3) and spans the 4-dimensional $\frac{\text{SU(3)}\times\text{Z}_{2}}{\text{SU(2)}\times\text{U(1)}}$ manifold, which is nonlinearly embedded in $\mathbb{R}^{8}$. SU(3) cannot be isomorphic or locally isomorphic to any other classical linear groups, which is in sharp contrast to the cases of even $N_{\text{f}}$ such as SU(2) and SU(4). Therefore, the new SU(3) antiferromagnetic order parameter we have discovered does not have a corresponding O($N_{\text{b}}$) order parameter. We temporarily refer to this new universality class as the chiral SU(3) universality class. As shown in the comparison in Table S4 summarizes the differences of this new universality class from the previously known ones in terms of symmetry and vacuum manifold. In addition to the symmetry analysis, our QMC results also provide an unbiased numerical determination of the critical exponents of this new universality class. In Table S5, we present a comparison of its critical exponents with those of the chiral Ising, chiral XY, and chiral Heisenberg universality classes, which more conclusively rules out the possibility that this new universality class belongs to those existing universality classes.
@@
Table S5: Comparison of critical exponents for different Gross-Neveu-Yukawa universality classes in $d=2+1$ with $N_{\mathrm{f}}=6$. The first row, chiral SU(3) denotes the universality class for SU(3) Dirac fermion Hubbard model, with exponents determined from nonequilibrium PQMC.
@@
chiral SU(3) (this work) 0.68(5) 0.55(5) 0.15(3)
@@
Figure S18: Schematic illustration of the relationship between O(NN) universality classes and Gross-Neveu universality classes. The O(NN) universality classes, including Ising, Heisenberg, and XY, describe spontaneous symmetry breaking with an O($N$) order parameter. The Gross-Neveu universality classes describe the situation where, along with spontaneous symmetry breaking of the bosonic order parameter, massless Dirac fermions also spontaneously acquire a mass. Within the Gross-Neveu universality classes, the traditional chiral O($N$) universality classes correspond one-to-one to the O($N$) universality classes, whereas the chiral SU(3) universality class in the Gross-Neveu universality classes does not possess an O($N$) order-parameter structure and instead represents a fermion-intrinsic universality class.
The nontrivial new Gross-Neveu universality class we discovered represents the first minimal example of a Gross-Neveu transition without a classical O($N$) correspondence. As illustrated in Fig. S18, our results demonstrate that the Gross-Neveu universality class encompasses a broader and nontrivial set with intrinsic fermionic characteristics. It is highly worthwhile in the future to further investigate the low-energy effective field theory of this phase transition. Since the bosonic field does not possess O($N_{\text{b}}$) symmetry, its form is likely different from the previous expression in Eq. (S35). We conjecture that it may take the following form:
$\mathcal{L}=\text{tr}\left(\bar{\psi}\not{\partial}\psi\right)+g\,\text{tr}\left(\bar{\psi}\Phi\psi\right)+\text{tr}\left[\left(\partial_{\mu}\Phi\right)^{2}\right]+r\,\text{tr}\left(\Phi^{2}\right)+\lambda\,\left[\text{tr}\left(\Phi^{2}\right)\right]^{2}+a\,\left[\text{tr}\left(\Phi^{3}\right)\right]^{2}+c\,\left[\text{tr}\left(\Phi^{2}\right)\right]^{4}\,$, (S37)
where $\Phi=\sum_{i=1}^{N_{\text{b}}}\phi_{i}\,\frac{\lambda_{i}}{2},N_{\text{b}}=8$ spans the 8-dimensional adjoint representation of SU(3). Compared with Eq. (S35), the marginal term $a\,\left[\text{tr}\left(\Phi^{3}\right)\right]^{2}$ is introduced to reduce the SO(8) symmetry of the bosonic field to $\text{SU(3)}\times\text{Z}_{2}$ symmetry, while the irrelevant term $c\,\left[\text{tr}\left(\Phi^{2}\right)\right]^{4}$ with $c>0$ is also introduced to ensure vacuum stability. Due to the SU(3) algebra identity $\text{tr}(\Phi^{3})=3\det\Phi$, only the $\phi_{8}$ component of the order parameter along the $\lambda_{8}$ direction with full rank contributes a nonzero value to $\left[\text{tr}\left(\Phi^{3}\right)\right]^{2}$. Therefore, when $a<0$, the SU(3)-symmetric bosonic field tends to spontaneously break toward the $\lambda_{8}$ direction. This is very different from the O($N$) symmetry breaking triggered by $r<0$ in the $\phi^{4}$ theory. Whether such a Gross-Neveu-Yukawa theory can capture the new universality class we discovered (for example, yielding critical exponents consistent with our numerical results) remains an open question for future study. Additionally, such an order parameter that carries an SU(3) bilinear representation rather than an O($N_{\text{b}}$) linear representation may also be constructed through a recent theoretical framework with tensorial criticality [90, 91, 92], which is also worth exploring.
S4.5 Sign problem behaviors and computational efficiency
@@
For this model, we set the parameters at the critical point $U=2.0(3)$, $\phi=0.1\pi$. The average sign as a function of imaginary time $\tau$ and size $L$ is shown in Fig. S19a. The sign problem for this model is significantly more severe than for the previous two models. For $\tau=0.25L^{z}$, the average sign is approximately $10^{-1}\sim 10^{-2}$, meaning that compared to the sign-problem-free case, $10$ to $100$ times more computational resources are needed to obtain reliable results. From Figs. S15 and S16 and Fig. 3 in the main text, it can be seen that $\tau=0.25L^{z}$ is in the nonequilibrium scaling region controlled by the critical point. Even though the ground state is not reached, its nonequilibrium scaling still reflects the quantum criticality of the ground state. In equilibrium QMC studies, it is typically necessary to set the imaginary time $\tau$ to more than $1.5$ times the size $L^{z}$ to reach the ground state. However, evolving for such a long time, the average sign decays to approximately $10^{-5}\sim 10^{-6}$. In Fig. S19b, we compare the average sign for $\tau=0.25L^{z}$ and $\tau=1.5L^{z}$. Since the computational error (according to Eq. S8) is inversely proportional to the average sign $\braket{\mathrm{sign}}$ (according to Eq. S8), we can measure the difference in computational efficiency by multiplying the ratio of the average sign $\braket{\mathrm{sign}}$ by the length of imaginary-time evolution, as follows:
$\mathrm{Efficiency~acceleration~factor}=\left(\frac{{1/\braket{\mathrm{sign}}_{\mathrm{eq.}}}}{{1/\braket{\mathrm{sign}}_{\mathrm{neq.}}}}\right)^{2}\frac{\tau_{\mathrm{eq.}}}{\tau_{\mathrm{neq.}}},$ (S38)